### v2

@@ 4 Trade-off rule
Exceeding the budget constraint in the population might be desirable if it achieves higher welfare and does not cost too much. How to model such trade-off depends on the specific welfare program. Section 4.1 first explores possible forms of the trade-off that are of practical relevance, and then focuses on the trade-off where the marginal cost per unit of violating the constraint has a known upper bound. The upper bound is context specific and as an example, Section 6 discusses how to set it for Medicaid expansion. Relaxing the constraint results in weakly higher welfare than that of the constrained optimal policy $g_{P}^{\ast}$. Section 4.2 derives a statistical rule that implements such trade-off in the sample, and is shown to be uniformly asymptotically welfare-efficient.
@@ 4.1 Form of the trade-off
Note that the original constrained optimization problem (2) may be reformulated as
$\max_{g\in\mathcal{G}}\inf_{\lambda\geq 0}\ W(g;P)-\lambda\cdot(B(g;P)-k).$ (15)
where $\lambda$ measures the marginal gain of relaxing the constraint. In many settings where the benefit is monetary, it may be natural to think for every dollar spent right above the budget constraint, there is a welfare gain of $\lambda$ dollars.
This formulation implies that policymakers are willing to enforce the constraint at all costs since $\lambda$ is unbounded. This formulation may not reflect the real objective, however, as policymakers might be willing to trade off violations of the constraint against gains in welfare only to a certain extent, bounding $\lambda\in[0,\overline{\lambda}]$ and resulting in a new objective function
$\max_{g\in\mathcal{G}}\min_{\lambda\in[0,\overline{\lambda}]}\ W(g;P)-\lambda\cdot(B(g;P)-k).$ (16)
Using the notation $(x)_{+}$ to denote the positive part of $x\in\mathbb{R}$, the inner optimization problem of (16) can be written as
$\min_{\lambda\in[0,\overline{\lambda}]}\ W(g;P)-\lambda\cdot(B(g;P)-k)=W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}.$
Note, however, this reformulated objective does not involve optimization over $\lambda$: the solution to (16) is equivalent to the solution to a non-smooth but piecewise linear optimization problem:
$\widetilde{g}_{P}\in\arg\max_{g\in\mathcal{G}}\ W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}.$ (17)
Note that this trade-off problem (17) is equivalent to an objective function faced by policymakers when they can finance the deficit $(B(g;P)-k)_{+}$ through borrowing at a fixed interest rate $\overline{\lambda}$, after converting welfare into monetary terms. Therefore, the solution $\widetilde{g}_{P}$ can relax the budget constraint and achieves weakly higher welfare than the constrained optimal policy $g_{P}^{\ast}$ for any data distribution $P$. The next lemma formalizes this observation.
@@
For any $\overline{\lambda}\in[0,\infty)$, the solution to the trade-off problem (17) achieves weakly higher welfare than the constrained optimum: $W(\widetilde{g}_{P};P)\geq W(\widetilde{g}_{P};P)-\overline{\lambda}\cdot(B(\widetilde{g}_{P};P)-k)_{+}\geq W(g_{P}^{\ast};P)$ for any distribution $P$. Moreover, the violation to the budget constraint is upper bounded by $\frac{W(\widetilde{g}_{P};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}$.
@@ 4.2 Welfare efficiency of the trade-off rule
Given a new objective function (17) that trades off the gain and the cost from violating the constraint, the goal is to derive a statistical rule that is likely to select eligibility policies that maximize the new objective function. Consider the trade-off statistical rule defined as
$\widehat{g}_{\text{tradeoff}}\in\arg\max_{g\in\mathcal{G}}\ \widehat{W}_{n}(g)-\overline{\lambda}\cdot(\widehat{B}_{n}(g)-k)_{+},$ (18)
where the subscript “tradeoff” highlights that this statistical rule is able to relax the constraint by trading off the gain and the cost from violating the constraint. Since $\widehat{g}_{\text{tradeoff}}$ solves a sample-analog version of (17), I show it consistently achieves the same value of (17) under weak conditions. Theorem 4.1 further verifies that it consistently achieves welfare that is weakly higher than $W(g_{P}^{\ast};P)$. I leave to future research to verify whether the trade-off rule is minimax rate optimal.
@@
Suppose Assumption 3.1 holds for the class of data distributions $\mathcal{P}$. Then the trade-off rule $\widehat{g}_{\text{tradeoff}}$ defined in (18) consistently achieves welfare that is weakly higher than $W(g_{P}^{\ast};P)$ for all $P\in\mathcal{P}$ and is uniformly asymptotically welfare-efficient under $\mathcal{P}$. Moreover, with probability approaching one, the violation to the budget constraint is upper bounded by $\frac{W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}$ uniformly over $\mathcal{P}$.
@@
To gain intuition for the above results, I note that under Assumption 3.1, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ consistently achieves the same value of (17) as $\widetilde{g}_{P}$ over repeated sample draws as $n\rightarrow\infty$ uniformly over $\mathcal{P}$. Furthermore, Lemma 4.1 shows the trade-off solution $\widetilde{g}_{P}$ achieves weakly higher welfare than the constrained optimal policy $g_{P}^{\ast}$ for any data distribution $P$. Therefore, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ is asymptotically welfare-efficient uniformly over $\mathcal{P}$. At the same time, larger trade-off coefficient $\bar{\lambda}$ implies smaller violation to the budget constraint, relative to the welfare gain $W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)$.
@@ 5 Monte Carlo Simulations
To ensure the practical relevance of the simulation, I calibrate to the distribution of the data from the OHIE, which is also used for empirical illustration in Section 6, based on Example 2.1 of Section 2. The benefit is defined to be the increase in the probability of reporting good subjective health after receiving Medicaid eligibility, and the cost is defined to be health care expenditure that needs to be reimbursed by Medicaid, in excess of the current policy. I defer the details on the construction of the estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$ to Section 6.1. Table 6.1 presents the basic summary statistics for the estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$. For the purpose of this simulation study, the OHIE represents the population $P$, and I take these estimates as the true benefit and cost $(\Gamma,R)$. Under this simulation design, I can solve for the constrained optimal policy as
$g_{P}^{\ast}\in\arg\max_{g\in\mathcal{G},\ B(g;P)\leq 0}W(g;P)$
where $\left(W(g;P),B(g;P)\right)$ are the sample analogs in the OHIE sample. The policy class $\mathcal{G}$ includes income thresholds that can vary with the number of children. Under the current definition of the cost, a policy is feasible if it incurs a negative cost. Table 6.1 shows the policy of expanding eligibility to the full population, while increasing health by 4%, violates the budget constraint by \$358 per enrollee. The constrained optimal policy therefore only makes a share of the population eligible. The maximum feasible welfare is given by $W(g_{P}^{\ast};P)=3.8\%$, an increase of 3.8% in reporting good subjective health. The cost associated with the constrained optimal policy is $B(g_{P}^{\ast};P)=-\$0.6$, meaning per enrollee $g_{P}^{\ast}$ costs \$0.6 less than the current policy.

### v3

@@ 4 Trade-off rule
Up to this point, the discussion has assumed that any budget violation is undesirable. If policymakers are willing to borrow, potentially at some penalty, to exceed the budget constraint in order to achieve higher welfare, then alternative rules might be preferable. Section 4.1 first formalizes this setting as a trade-off problem. Section 4.2 derives a statistical rule that implements such trade-off in the sample, and is shown to be uniformly asymptotically welfare-efficient.
@@ 4.1 Form of the trade-off
Exceeding the budget can have negative economic consequences, such as increased borrowing costs or reduced funding for other programs. However, it may not be severe enough to entirely outweigh the perceived welfare gains. Consider a new objective function in which the policymaker, while operating within the budget constraint, seeks to allocate resources efficiently to maximize welfare, but once the budget is exceeded, must trade off the welfare gains from additional spending against the penalty associated with the overrun. Let $\overline{\lambda}>0$ denote this penalty. Using the notation $(x)_{+}$ to represent the positive part of $x\in\mathbb{R}$, the objective function can be written as [^8] This objective function can also be motivated as policymakers might be willing to trade off violations of the constraint against gains in welfare only to a certain extent, bounding the marginal gain of relaxing the constraint $\lambda\in[0,\overline{\lambda}]:$ $\max_{g\in\mathcal{G}}\min_{\lambda\in[0,\overline{\lambda}]}\ W(g;P)-\frac{\lambda}{r}\cdot(B(g;P)-k).$
$\max_{g\in\mathcal{G}}V(g;P)\text{ where }V(g;P)\coloneqq W(g;P)-\frac{\overline{\lambda}}{r}\cdot(B(g;P)-k)_{+}.$ (9)
Here $r>0$ denotes the rate at which monetary units are converted into welfare units, since welfare is often not measured in monetary terms whereas the budget is.
This objective function (9) is a non-smooth but piecewise linear optimization problem. Denote its solution to be:
$\widetilde{g}_{P}\in\arg\max_{g\in\mathcal{G}}\ V(g;P),$ (10)
and note that the solution $\widetilde{g}_{P}$ can relax the budget constraint and achieves weakly higher welfare than the constrained optimal policy $g_{P}^{\ast}$ for any data distribution $P$. The next lemma formalizes this observation.
@@
For any $\overline{\lambda}\in[0,\infty)$, the solution to the trade-off problem (10) achieves weakly higher welfare than the constrained optimum: $W(\widetilde{g}_{P};P)\geq V(\widetilde{g}_{P};P)\geq W(g_{P}^{\ast};P)$ for any distribution $P$. Moreover, the violation to the budget constraint is upper bounded by $\frac{r\cdot(W(\widetilde{g}_{P};P)-W(g_{P}^{\ast};P))}{\overline{\lambda}}$.
@@ 4.2 Welfare efficiency of the trade-off rule
Given a new objective function (10) that trades off the gain and the cost from violating the constraint, the goal is to derive a statistical rule that is likely to select eligibility policies that maximize the new objective function $V(g;P)$. Consider the trade-off statistical rule defined as
$\widehat{g}_{\text{tradeoff}}\in\arg\max_{g\in\mathcal{G}}\ \widehat{V}_{n}(g)\text{ where }\widehat{V}_{n}(g)\coloneqq\widehat{W}_{n}(g)-\frac{\overline{\lambda}}{r}\cdot(\widehat{B}_{n}(g)-k)_{+},$ (11)
where the subscript “tradeoff” highlights that this statistical rule is able to relax the constraint by trading off the gain and the cost from violating the constraint. Since $\widehat{g}_{\text{tradeoff}}$ solves a sample-analog version of (10), I show it consistently achieves the maximal value of $V(g;P)$ under weak conditions in Lemma 4.2. Theorem 4.1 further verifies that it consistently achieves welfare that is weakly higher than $W(g_{P}^{\ast};P)$. I leave to future research to verify whether the trade-off rule is minimax rate optimal.
Lemma 4.2.
Suppose Assumption 3.1 holds for the class of data distributions $\mathcal{P}$. Then the trade-off rule $\widehat{g}_{\text{tradeoff}}$ defined in (11) consistently achieves the maximized value of $V(g;P)$, namely $V(\widetilde{g}_{P};P).$
@@
For the class of data distributions $\mathcal{P}$ satisfying Assumption 3.1, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ consistently achieves welfare that is weakly higher than $W(g_{P}^{\ast};P)$ for all $P\in\mathcal{P}$ and is therefore uniformly asymptotically welfare-efficient under $\mathcal{P}$. Moreover, with probability approaching one, the violation to the budget constraint is upper bounded by $\frac{r\cdot(W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P))}{\overline{\lambda}}$ uniformly over $\mathcal{P}$.
@@
To gain intuition for the above results, note that Lemma 4.2 shows the trade-off rule $\widehat{g}_{\text{tradeoff}}$ uniformly consistently achieves $V(\widetilde{g}_{P};P),$ which by Lemma 4.1 is weakly higher than the welfare achieved by the constrained optimal policy $g_{P}^{\ast}$ for any data distribution $P$. Therefore, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ is asymptotically welfare-efficient uniformly over $\mathcal{P}$. At the same time, larger $\bar{\lambda}$ implies smaller violation to the budget constraint, relative to the welfare gain $W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)$.
4.3 Medicaid expansion: empirical illustration
Example 2.1 of Section 2 explains a more flexible Medicaid expansion policy that would allow the income thresholds to vary with the number of children. In this example, while the budget constraint is set equal to the cost of the current policy, policymakers would be willing to exceed the budget constraint, potentially at some penalty, in order to achieve higher welfare. This subsection uses this example, together with data from the Oregon Medicaid Health Insurance Experiment (OHIE), to illustrate the trade-off rule and compare it with the sample-analog rule and its modification.
4.3.1 Data
The original OHIE sample consists of 74,922 individuals (representing 66,385 households). Of these, 26,423 individuals responded to the initial mail survey, which collects information on income as percentage of the federal poverty level and number of children, which are the characteristics of interest for targeting ($X_{i}$). [^9] More accurately, I follow Sacarny et al. (2020) to approximate number of children by the number of family members under age 19 living in house as reported on the initial mail survey. I exclude individuals who did not respond to the initial survey from my sample, which differs from the sample analyzed in Finkelstein et al. (2012) as I focus on individuals who responded both to the initial and the main surveys from the OHIE. Due to this difference, the expansion policies selected using my sample do not directly carry their properties to the population underlying the original OHIE sample, as the distributions of $X$ differ. After one year, the main survey collects data related to health ($Y_{i}$), health care utilization ($H_{i}$) and actual enrollment in Medicaid ($M_{i}$), which allows me to construct estimates for the benefit and cost of Medicaid eligibility $(\tau,C)$. Therefore I further exclude individuals who did not respond to the main survey from my sample.
For health ($Y_{i}$), I follow the binary measurement in Finkelstein et al. (2012) based on self-reported health, where an answer of “poor/fair” is coded as $Y_{i}=0$ and “excellent/very good/good” is coded as $Y_{i}=1$. For health care utilization ($H_{i}$), the study collected measures of utilization of prescription drugs, outpatient visits, ER visits, and inpatient hospital visits. Finkelstein et al. (2012) annualize these utilization measures to turn these into spending estimates, weighting each type by its average cost (expenditures inflated with the CPI-U to 2007 dollars) among low-income publicly insured non-elderly adults in the Medical Expenditure Survey (MEPS). Note that health and health care utilization are not measured at the same scale, which requires rescaling when I consider the trade-off between the two. I address this issue in Section 4.3.2. Lastly, since the enrollment in Medicaid still requires an application, not everyone eligible in the OHIE eventually enrolled in Medicaid, which implies $M_{i}\leq D_{i}$.
Given the setup of the OHIE, Medicaid eligibility ($D_{i}$) is random conditional on household size (number of adults in the household) entered on the lottery sign-up form and survey wave. While the original experimental setup would ensure randomization given household size, the OHIE had to adjust randomization for later waves of survey respondents (see the Appendix of Finkelstein et al. (2012) for more details). Denote the confounders (household size and survey wave) with $V_{i}$, and define the propensity score as $p(V_{i})=\operatorname{Pr}\{D_{i}=1\mid V_{i}\}$. If the propensity score is known, then the construction of the estimates follows directly from the formula (5). However, the adjustment for later survey waves means I need to estimate the propensity score, and I adapt the formula (5) following Athey and Wager (2021) to account for the estimated propensity score.
Specifically, define the conditional expectation function (CEF) of a random variable $U_{i}$ as $\gamma^{U}=\mathbb{E}[U_{i}\mid V_{i},D_{i}$]. Since $V_{i}$ in my case is discrete, I use a fully saturated model to estimate the propensity score $\widehat{p}(V_{i})$ and the CEF $\widehat{\gamma}^{U}(V_{i},D_{i})$. I then form the estimated Horvitz-Thompson weight with the estimated propensity score as $\widehat{\alpha}(V_{i},D_{i})=\frac{D_{i}}{\widehat{p}(V_{i})}-\frac{1-D_{i}}{1-\widehat{p}(V_{i})}$. For health benefit due to Medicaid eligibility, define the estimate $\tau_{i}^{\ast}=\widehat{\gamma}^{Y}(V_{i},1)-\widehat{\gamma}^{Y}(V_{i},0)+\widehat{\alpha}(V_{i},D_{i})\cdot\left(Y_{i}-\widehat{\gamma}^{Y}(V_{i},D_{i})\right)$. For the cost due to Medicaid eligibility, define the estimate $C_{i}^{\ast}=\widehat{\gamma}^{Z}(V_{i},1)+\frac{D_{i}}{\widehat{p}(W_{i})}\cdot\left(Z_{i}-\widehat{\gamma}^{Z}(V_{i},D_{i})\right)$ where $Z_{i}=M_{i}\cdot H_{i}$. Since an eligible individual only incurs cost to Medicaid if enrolled, I need to account for imperfect take-up in forming $C_{i}^{\ast}$.
Table 4.1: Summary statistics of the OHIE sample by number of children
Number of children Sample size Sample mean of $\tau_{i}^{\ast}$ Sample mean of $C_{i}^{\ast}$
0 5,758 3.16% \$1,974
(0.01) (110)
1 1,736 10.34% \$1,615
(0.02) (195)
$\geq 2$ 2,641 1.55% \$1,451
(0.02) (129)
Full sample 10,135 3.97% \$1,776
(0.01) (79)
Notes: This table presents summary statistics on the sample of individuals who responded to both the initial and the main surveys from the Oregon Health Insurance Experiment (the OHIE sample). The first three rows represent individuals living with different number of children (family members under age 19), and the last row is the aggregate. The estimate for benefit $\tau_{i}^{\ast}$ is an estimate for the increase in the probability of an individual reporting “excellent/very good/good” on self-reported health (as opposed to “poor/fair”) after receiving Medicaid eligibility. The estimate for cost $C_{i}^{\ast}$ is an estimate for individual’s health care expenditure that needs to be reimbursed by Medicaid. Standard errors are shown in parentheses below.
Table 4.1 presents the summary statistics. While Appendix B.3 argues the estimation errors in $\tau_{i}^{\ast}$ and $C_{i}^{\ast}$ are asymptotically negligible, in finite samples, the cost estimates are highly variable, resulting in noisy estimate $\widehat{B}_{n}(g)$.
To formalize the budget constraint requiring that the average cost of any proposed policy does not exceed that of the status quo 2014 Medicaid expansion, I calibrate the per capita cost under the 2014 policy. Following Finkelstein et al. (2012), who cite Wallace et al. (2008), Medicaid spending among individuals comparable to the Oregon Health Insurance Experiment (OHIE) participants was approximately \$3,000 per enrollee in Oregon in 2004, which corresponds to about \$3,300 in 2007 dollars. Under the 2014 policy, I adjust for imperfect take-up by defining the per capita cost as
$k=\mathbb{E}_{P}[M(1)g_{2014}(X)]\cdot\mathdollar 3,300=\mathdollar 1,377$ (12)
where $M(1)$ denotes enrollment status if offered Medicaid eligibility and $g_{2014}(x)=\mathbf{1}\{\text{income}\leq 138\%\}$ represents the status quo 2014 expansion policy of providing eligibility to all adults with income up to 138%. The enrollment rate under the status quo 2014 policy $\mathbb{E}_{P}[M(1)g_{2014}(X)]$ is based on point estimate from the OHIE. [^10] The theoretical results in this paper are developed for a fixed $k$. Addressing the estimation error in $k$ is left to future work.
4.3.2 Budget-constrained Medicaid expansion
Figure 4.1 summarizes the selected expansion policies, which are income thresholds specific to the number of children. The sample-analog rule $\widehat{g}_{\text{sample}}$ chooses to restrict Medicaid eligibility, especially lowering the income threshold for childless individuals far below the current level, and the estimated welfare is 3.79% increase in reporting good subjective health. The budget estimate for the selected policy is \$1,311, slightly below the threshold of $k=\mathdollar 1,377$, as $\widehat{g}_{\text{sample}}$ imposes the sample-analog version of the budget constraint. However, due to the large variation in the cost estimates as illustrated in Table 4.1, meeting the sample budget constraint still involves uncertainty about whether the selected policy meets the budget constraint in the population as argued in Proposition 3.1. After taking into account of estimation uncertainty, the modified sample-analog rule $\widehat{g}_{\alpha=5\%}$ is much more conservative than the sample-analog rule $\widehat{g}_{\text{sample}}$. The budget estimate of the selected policy \$1,174, with a standard error of 66, making it statistically significantly below the budget constraint at the conventional 5% level. The welfare estimate of the selected policy is lower at 3.43%. One can reduce the conservativeness by increasing $\alpha$ and therefore lowering the statistical guarantee that the selected policy meets the budget constraint in the population. I examine the results under higher values of $\alpha$ in Appendix D.1.
To construct the trade-off rule $\widehat{g}_{\text{tradeoff}}$ as proposed in Section 4.1, I need to specify both the penalty parameter, $\overline{\lambda}$ and the conversion rate, $r$. For illustration, I assume that exceeding the budget incurs a full repayment of the overrun, implying $\overline{\lambda}=1.$ In my empirical illustration, the budget constraint is in terms of monetary value. The objective function, however, is measured based on self-reported health, which does not directly translate to a monetary value. Finkelstein et al. (2019) converts self-reported health into value of a statistical life year (VSLY) based on existing estimates. Specifically, a conservative measure for the increase in quality-adjusted life year (QALY) when self-reported health increases from “poor/fair” to “excellent/very good/good” is roughly 0.6. The “consensus” estimate for the VSLY for one unit of QALY from Cutler (2004) is \$100,000 for the general US population. I therefore follow Finkelstein et al. (2019) and set $r=(0.6\cdot\mathdollar 100,000)=\mathdollar 60,000.$ The trade-off objective function is therefore
$\max_{g\in\mathcal{G}}V(g;P)\text{ where }V(g;P)\coloneqq W(g;P)-\frac{1}{\mathdollar 60,000}\cdot(B(g;P)-k)_{+}.$ (13)
The trade-off rule $\widehat{g}_{\text{tradeoff}}$ based on (13) chooses to assign Medicaid eligibility to more individuals, and to raise the income thresholds above the current level for those with children. Therefore the welfare estimate for the selected policy is higher at 3.86% and budget estimate for the selected policy is \$1,402, slightly above the budget constraint. The higher level occurs because on average the benefit estimates are positive, and the trade-off rule finds that the additional health benefit from violating the budget constraint exceeds the cost of doing so, especially those with children. While a traditional cost-benefit analysis would also recommend prioritizing those with children based on Table 4.1, the trade-off rule offers a more interpretable recommendation by directly selecting income thresholds.
However, setting $\overline{\lambda}=1$ assumes that the penalty is limited to repaying the budget overrun, potentially understating the true penalty for policymakers, as it does not account for additional economic costs associated with exceeding the budget, such as reduced funding for other programs or lower quality of Medicaid services. Accordingly, in Appendix D.1, I explore how the results change when $\overline{\lambda}$ is increased, which is equivalent to making budget overruns more costly to the policymaker. For example, with $\overline{\lambda}=1.68$, the welfare and budget estimates for the selected policy essentially drop to those of $\widehat{g}_{\text{sample}}$.
Figure 4.1: More flexible Medicaid expansion policies
(a) policy selected by the sample-analog rule $\widehat{g}_{\text{sample}}$
(b) policy selected by the modified rule $\widehat{g}_{\alpha=5\%}$
(c) policy selected by the trade-off rule $\widehat{g}_{\text{tradeoff}}$
Notes: This figure plots the more flexible Medicaid expansion policies selected by statistical rules based on results from the OHIE. The horizontal dashed line marks the income thresholds under the current expansion policy, which is 138% regardless of the number of children in a household. The horizontal solid lines mark the more flexible policy selected by various statistical rules, i.e. income thresholds that can vary with number of children. For each number of children, I also plot the underlying income distribution to visualize individuals below the thresholds. Panel (a) plots the policy selected by the sample-analog rule $\widehat{g}_{\text{sample}}$. Panel (b) plots the policy selected by the modified sample-analog rule $\widehat{g}_{\alpha=5\%}$. Panel (c) plots the policy selected by the trade-off rule $\widehat{g}_{\text{tradeoff}}$.
In this example, the policies chosen by different rules differ substantially. A natural question is how policymakers should choose the rules. In this Medicaid example, it is reasonable to assume that policymakers can borrow, potentially at some penalty, to exceed the budget constraint in order to achieve higher welfare. Therefore the trade-off rule $\widehat{g}_{\text{tradeoff}}$ is theoretically attractive. It selects policies that achieve at least the maximum feasible welfare, while accounting for the penalty $\overline{\lambda}$ from violating the budget constraint. Because the amount of budget violations depend on the choice of $\overline{\lambda}$ as shown in Theorem 4.1, one can assess this trade-off for their particular settings by varying $\overline{\lambda}$. However, in other context where policymakers are financially conservative and even minor violation to the budget constraint is unacceptable, then the modified sample-analog rule $\widehat{g}_{\alpha}$, with a pre-specified significance level $\alpha$, is theoretically attractive. It offers a statistical guarantee that the budget constraint will be met in the population with high confidence. However, the selected eligibility policy may appear very conservative, as illustrated in the Medicaid expansion example above.
@@ 5 Monte Carlo Simulations
To ensure the practical relevance of the simulation, I calibrate to the distribution of the data from the OHIE, which is also used for empirical illustration in Section 4.3, based on Example 2.1 of Section 2.
For the purpose of this simulation study, the OHIE represents the population $P$, and I take the estimates $(\tau_{i}^{\ast},C_{i}^{\ast})$ constructed in Section 4.3.1 as the true benefit and cost $(\tau,C)$. Under this simulation design, I can solve for the constrained optimal policy as
$g_{P}^{\ast}\in\arg\max_{g\in\mathcal{\widetilde{G}},\ B(g;P)\leq k}W(g;P)$
where $\left(W(g;P),B(g;P)\right)$ are the sample analogs in the OHIE sample. The policy class $\mathcal{\widetilde{G}}$ includes income thresholds that can vary with the number of children as described in Equation (19) of Appendix D.2, which is slightly coarsened than the class used in Section 4.3.2. The maximum feasible welfare is given by $W(g_{P}^{\ast};P)=3.76\%$, an increase of 3.76% in reporting good subjective health. The cost associated with the constrained optimal policy is $B(g_{P}^{\ast};P)=\mathdollar 1,340$, slightly below the constraint $k=\mathdollar 1,377$ as defined in Equation (12).